# Matryoshka: Optimization of Dynamic Diverse Quantum Chemistry Systems via Elastic Parallelism Transformation

### v1

@@ 1. Introduction
We conduct a comprehensive evaluation of Matryoshka on 13 representative QC systems (Chignolin, DNA, Crambin, etc.), demonstrating a remarkable speedup of up to 5.9x compared to state-of-the-art solutions. Notably, existing state-of-the-art approaches, while preserving original accuracy, have not surpassed the simulation scale of more than 1,000 atoms (Mironov et al., 2019; Sun, 2015; Sun et al., 2018, 2020; Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023). Matryoshka achieves a breakthrough, maintaining original ab initio accuracy while simulating 11,259 atoms using a single GPU in one day (19.5 hrs).
@@ 2.3. State-of-the-art
Recent efforts to accelerate research on large-scale QC systems with ab initio accuracy have pursued two primary approaches (Mironov et al., 2019; Ufimtsev and Martínez, 2008a, b; Ufimtsev and Martinez, 2009a, b; Asadchev et al., 2010; Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023; Barca et al., 2020). One approach involves utilizing a greater number of CPU cores for distributed computing acceleration, a solution that consumes vast computational resources and is prohibitively expensive (Mironov et al., 2019).The other focuses on efficiently porting QC systems to GPU architecture using High Performance Computing (HPC) techniques. However, as discussed earlier, the current approaches face significant challenges due to dynamic diversity, which considerably limits the potential of GPUs(Ufimtsev and Martínez, 2008a, b; Ufimtsev and Martinez, 2009a, b; Asadchev et al., 2010; Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023; Barca et al., 2020). To the best of our knowledge, the industry-recognized state-of-the-art approaches maintaining original ab initio accuracy have yet to surpass the scale of simulating more than 1,000 atoms (Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023; Barca et al., 2020). Consequently, the design at the system level to efficiently support larger-scale QC systems with fewer computational resources emerges as a crucial and challenging endeavor.
@@ 9. Related Work
CPU-centric. Traditional CPU-based approaches (Almlöf et al., 1982; Pulay, 1980; Cremer and Gauss, 1986; Häser and Ahlrichs, 1989) primarily focus on minimizing computation and can be divided into two strategies: reducing the number of iterations (Almlöf et al., 1982; Cremer and Gauss, 1986) and limiting the number of integrals recalculated per iteration (Pulay, 1980; Cremer and Gauss, 1986). Although these improvements significantly reduce computation, they do not consider computer hardware, resulting in suboptimal computational efficiency. Recently, several studies (Mironov et al., 2019; Valiev et al., 2010; Alexeev et al., 2002, 2007; Fletcher et al., 2000; Liu et al., 2014; Chow et al., 2015; Alexeev et al., 2014; Shan et al., 2014, 2015) have concentrated on developing efficient implementations using distributed CPU nodes. Some works aim to design distributed parallel algorithms (Mironov et al., 2019; Valiev et al., 2010; Alexeev et al., 2002, 2007; Fletcher et al., 2000), while others address the prevalent load imbalance issues in the field (Liu et al., 2014; Chow et al., 2015; Alexeev et al., 2014; Shan et al., 2014, 2015). However, the limitations in computational capability of CPUs hinder their further advancement.
GPU-centric. Yasudo (Yasuda, 2008) was among the first to implement GPU technology for the evaluation of ERIs, resulting in substantial speed improvements when compared to CPU-based computations. Following this, Ufimtsev and Martinez (Ufimtsev and Martínez, 2008a, b; Ufimtsev and Martinez, 2009a, b) conducted a series of studies that implemented a full Fock build on GPUs, wherein each thread is mapped to a different fundamental integral class. However, their works mainly depended on single-precision arithmetic, leading to considerable computational errors. Asadchev et al. (Asadchev et al., 2010) introduced a Rys quadrature ERI method on GPUs that utilized double precision, but overlooked the data transfer costs between CPUs and GPUs. Miao et al. (Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023) and Barca et al. (Barca et al., 2020) employed the HGP method to reduce the computational expenses associated with the evaluation of ERIs, ultimately achieving remarkable speedups. Despite these advancements, these approaches did not account for the dynamic diversity inherent in QC systems, resulting in under-utilization.

### v2

@@ 1. Introduction
We conduct a comprehensive evaluation of Matryoshka on 13 representative QC systems (Chignolin, DNA, Crambin, etc.), demonstrating a remarkable speedup of up to 5.9x compared to state-of-the-art solutions. Notably, existing state-of-the-art approaches, while preserving original accuracy, have not surpassed the simulation scale of more than 1,000 atoms (Mironov et al., 2019a; Sun, 2015; Sun et al., 2018, 2020; Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023). Matryoshka achieves a breakthrough, maintaining original ab initio accuracy while simulating 11,259 atoms using a single GPU in one day (19.5 hrs).
@@ 2.3. State-of-the-art
Recent efforts to accelerate research on large-scale QC systems with ab initio accuracy have pursued two primary approaches (Mironov et al., 2019a; Ufimtsev and Martínez, 2008a, b; Ufimtsev and Martinez, 2009a, b; Asadchev et al., 2010; Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023; Barca et al., 2020). One approach involves utilizing a greater number of CPU cores for distributed computing acceleration, a solution that consumes vast computational resources and is prohibitively expensive (Mironov et al., 2019a).The other focuses on efficiently porting QC systems to GPU architecture using High Performance Computing (HPC) techniques. However, as discussed earlier, the current approaches face significant challenges due to dynamic diversity, which considerably limits the potential of GPUs(Ufimtsev and Martínez, 2008a, b; Ufimtsev and Martinez, 2009a, b; Asadchev et al., 2010; Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023; Barca et al., 2020). To the best of our knowledge, the industry-recognized state-of-the-art approaches maintaining original ab initio accuracy have yet to surpass the scale of simulating more than 1,000 atoms (Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023; Barca et al., 2020). Consequently, the design at the system level to efficiently support larger-scale QC systems with fewer computational resources emerges as a crucial and challenging endeavor.
@@ 9. Related Work
CPU-centric. Traditional CPU-based approaches (Almlöf et al., 1982; Pulay, 1980; Cremer and Gauss, 1986; Häser and Ahlrichs, 1989) primarily focus on minimizing computation and can be divided into two strategies: reducing the number of iterations (Almlöf et al., 1982; Cremer and Gauss, 1986) and limiting the number of integrals recalculated per iteration (Pulay, 1980; Cremer and Gauss, 1986). Although these improvements significantly reduce computation, they do not consider computer hardware, resulting in suboptimal computational efficiency. Recently, several studies (Mironov et al., 2019a; Valiev et al., 2010a; Alexeev et al., 2002, 2007; Fletcher et al., 2000; Liu et al., 2014; Chow et al., 2015; Alexeev et al., 2014; Shan et al., 2014, 2015) have concentrated on developing efficient implementations using distributed CPU nodes. Some works aim to design distributed parallel algorithms (Mironov et al., 2019a; Valiev et al., 2010a; Alexeev et al., 2002, 2007; Fletcher et al., 2000), while others address the prevalent load imbalance issues in the field (Liu et al., 2014; Chow et al., 2015; Alexeev et al., 2014; Shan et al., 2014, 2015). However, the limitations in computational capability of CPUs hinder their further advancement.
GPU-centric. The rapid advancements in general-purpose GPU technology have motivated many packages (Mironov et al., 2019b; Valiev et al., 2010b; Frisch et al., 2016; Kühne et al., 2020; Asadchev and Valeev, 2023a, b, 2024) to transition from traditional CPU to GPU architectures. Additionally, a variety of GPU-specific methods have emerged. Yasudo (Yasuda, 2008) was among the first to implement GPU technology for the evaluation of ERIs, resulting in substantial speed improvements when compared to CPU-based computations. Following this, Ufimtsev and Martinez (Ufimtsev and Martínez, 2008a, b; Ufimtsev and Martinez, 2009a, b) conducted a series of studies that implemented a full Fock build on GPUs, wherein each thread is mapped to a different fundamental integral class. However, their works mainly depended on single-precision arithmetic, leading to considerable computational errors. Asadchev et al. (Asadchev et al., 2010) introduced a Rys quadrature ERI method on GPUs that utilized double precision, but overlooked the data transfer costs between CPUs and GPUs. Miao et al. (Miao and Merz, 2013, 2015; Manathunga et al., 2021, 2023) and Barca et al. (Barca et al., 2020) employed the HGP method to reduce the computational expenses associated with the evaluation of ERIs, ultimately achieving remarkable speedups. Despite these advancements, these approaches did not account for the dynamic diversity inherent in QC systems, resulting in under-utilization.